a shop keeper sells two phones for rs 6000, one at a gain of 25 percent and the other at a loss of 25 percent. find the overall gain or loss percent.

Answer by Edwin McCravy(20060)

a shop keeper sells two phones for rs 6000, one at a gain of 25 percent and the other at a loss of 25 percent. find the overall gain or loss percent.
That is not enough information.

If the first phone sold for rs 1000 and the second phone at rs 5000, then
1. The gain of 25 percent of the first phone is (1000)(0.25) = rs 250.
2. The loss of 25 percent of the second phone is (5000)(-0.25) = rs - rs 1250.
3. That would be a total LOSS of rs 250 - rs 1250 = -rs 1000 a LOSS of rs 1250.
4.  That would be a total LOSS of 1250/6000 = 20.8 percent.

If the first phone sold for rs 2000 and the second phone at rs 4000, then
1. The gain of 25 percent of the first phone is (2000)(0.25) = rs 500.
2. The loss of 25 percent of the second phone is (4000)(-0.25) = - rs 1000.
3. That would be a total LOSS of rs 500 - rs 1000 = -rs 500 a LOSS of rs 500.
4.  That would be a total LOSS of 500/6000 = 8.3 percent.


If the first phone sold for rs 4000 and the second phone at rs 2000, then
1. The gain of 25 percent of the first phone is (4000)(0.25) = rs 1000.
2. The loss of 25 percent of the second phone is (2000)(-0.25) = - rs 500.
3. That would be a total GAIN of rs 1000- rs 500 = -rs 500 a GAIN of rs 500.
4.  That would be a total GAIN of 500/6000 = 8.3 percent.

If the first phone sold for rs 5000 and the second phone at rs 1000, then
1. The gain of 25 percent of the first phone is (5000)(0.25) = rs 1250.
2. The loss of 25 percent of the second phone is (1000)(-0.25) = rs 1250.
3. That would be a total GAIN of rs 1250 - rs 250 = rs 1000 a GAIN of rs 1250.
4.  That would be a total GAIN of 1250/6000 = 20.8 percent.

We need more information.

Edwin